# 3D-free meets 3D priors: Novel View Synthesis from a Single Image with Pretrained Diffusion Guidance

### v2

@@ 5 Experiments and Results
We utilize the HawkI-Syn Kothandaraman et al. (2023) and HawkI-Real Kothandaraman et al. (2023) datasets that feature complex scenes with multiple foreground objects and background. Both datasets provide images and text prompts to the model. Explain the subset you used for your results, how you created the, how many images per sample you used for evaluation, etc.
@@
We compare our method with state-of-the-art view synthesis methods: Zero123++ Shi et al. (2023a) and Stable Zero123 for 3D-based methods and HawkI Kothandaraman et al. (2023) for 3D-free method.
@@ 5.1 Qualitative analysis
Our model is able to generate distinct viewpoints in the camera control to ensure consistency across generated views. Details are mentioned in the Zero123++ Shi et al. (2023a). We present qualitative representative results and comparisons with Zero123++ Shi et al. (2023a), HawkI Kothandaraman et al. (2023), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 5 and 6. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 5 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, our model accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints.
@@ 5.2 Quantitative Evaluation
Following prior work Shi et al. (2023a); Kothandaraman et al. (2023); Liu et al. (2023a), we evaluate our method using six metrics - (i) LPIPS Zhang et al. (2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score Radford et al. (2021) and View-CLIP Score: Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{\text{elev}},\alpha_{\text{azi}})$ viewpoint, while the View-CLIP Score focuses specifically on the viewpoint. Higher values are preferred. (iii) CLIP-I Ruiz et al. (2023), CLIPD, and View-CLIPD: Evaluate image fidelity and directional similarity. CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. CLIPD [^1] https://huggingface.co/docs/diffusers/conceptual/evaluation and View-CLIPD assess the consistency of transformations between two images, using text prompts similar to CLIP-Score and View-CLIP scores. Higher values are preferred. Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 and Table 2 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how our models stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches.
@@
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ View-CLIP $\uparrow$ CLIPD $\uparrow$ View-CLIPD $\uparrow$ CLIP-I $\uparrow$
HawkI-Syn $(30^{\circ},30^{\circ})$ Ours 0.5667 29.9125 19.7100 0.0534 0.0568 0.8334
HawkI-Syn $(30^{\circ},30^{\circ})$ HawkI 0.5998 27.6516 18.6122 0.0405 0.0549 0.8108
HawkI-Syn $(30^{\circ},30^{\circ})$ Zero123++ 0.5793 29.8709 18.4740 0.0552 0.0540 0.8330
HawkI-Syn $(30^{\circ},30^{\circ})$ Stable Zero123 0.6853 20.5335 18.0746 0.0528 0.0520 0.6181
HawkI-Syn $(30^{\circ},270^{\circ})$ Ours 0.5791 29.2654 19.1836 0.0553 0.0611 0.8274
HawkI-Syn $(30^{\circ},270^{\circ})$ HawkI 0.5996 29.0022 19.1088 0.0321 0.0498 0.8438
HawkI-Syn $(30^{\circ},270^{\circ})$ Zero123++ 0.5911 28.9917 19.1491 0.0482 0.0597 0.7917
HawkI-Syn $(30^{\circ},270^{\circ})$ Stable Zero123 0.6441 21.2231 18.3738 0.0406 0.0403 0.6152
HawkI-Real $(30^{\circ},30^{\circ})$ Ours 0.6075 30.9333 20.7728 0.0829 0.0730 0.8333
HawkI-Real $(30^{\circ},30^{\circ})$ HawkI 0.6405 28.5431 18.5430 0.0506 0.0468 0.8302
HawkI-Real $(30^{\circ},30^{\circ})$ Zero123++ 0.6679 29.8661 19.9530 0.0624 0.0467 0.8049
HawkI-Real $(30^{\circ},30^{\circ})$ Stable Zero123 0.6624 22.8604 18.7536 0.0523 0.0658 0.6740
HawkI-Real $(30^{\circ},270^{\circ})$ Ours 0.6180 30.0677 19.6756 0.0652 0.0709 0.8350
HawkI-Real $(30^{\circ},270^{\circ})$ HawkI 0.6497 28.4452 18.4231 0.0574 0.0577 0.8236
HawkI-Real $(30^{\circ},270^{\circ})$ Zero123++ 0.6935 27.2773 19.5926 0.0605 0.0698 0.7212
HawkI-Real $(30^{\circ},270^{\circ})$ Stable Zero123 0.6466 23.6915 17.8817 0.0593 0.0658 0.6726
@@
Table 1: Quantitative Results. Evaluation on six metrics demonstrate the superior results of our method over prior work.
@@
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ View-CLIP $\uparrow$ CLIPD $\uparrow$ View-CLIPD $\uparrow$ CLIP-I $\uparrow$
HawkI-Syn $(-20^{\circ},210^{\circ})$ Ours 0.5521 28.5599 17.9664 0.0502 0.0565 0.8535
HawkI-Syn $(-20^{\circ},210^{\circ})$ HawkI 0.5747 27.0915 16.9330 0.0403 0.0477 0.8370
HawkI-Syn $(-20^{\circ},210^{\circ})$ Zero123++ 0.7193 23.9836 16.7291 0.0496 0.0886 0.7389
HawkI-Syn $(-20^{\circ},210^{\circ})$ Stable Zero123 0.6938 20.3084 16.8253 0.0403 0.0400 0.5826
HawkI-Syn $(-20^{\circ},330^{\circ})$ Ours 0.5672 29.1569 18.1494 0.0556 0.0535 0.8656
HawkI-Syn $(-20^{\circ},330^{\circ})$ HawkI 0.5967 27.6743 17.3102 0.0520 0.0529 0.8409
HawkI-Syn $(-20^{\circ},330^{\circ})$ Zero123++ 0.6662 24.6859 17.5781 0.0688 0.0962 0.7383
HawkI-Syn $(-20^{\circ},330^{\circ})$ Stable Zero123 0.6201 22.3172 16.2088 0.0499 0.0542 0.7034
HawkI-Real $(-20^{\circ},210^{\circ})$ Ours 0.6222 29.6425 18.8914 0.0744 0.0709 0.7868
HawkI-Real $(-20^{\circ},210^{\circ})$ HawkI 0.6364 27.3844 17.5037 0.0568 0.0507 0.7821
HawkI-Real $(-20^{\circ},210^{\circ})$ Zero123++ 0.7739 24.3284 18.7046 0.0713 0.0951 0.6798
HawkI-Real $(-20^{\circ},210^{\circ})$ Stable Zero123 0.6637 21.2952 17.2609 0.0556 0.0704 0.5829
HawkI-Real $(-20^{\circ},330^{\circ})$ Ours 0.6135 30.6951 18.9849 0.0506 0.0371 0.8500
HawkI-Real $(-20^{\circ},330^{\circ})$ HawkI 0.6500 28.5102 17.4244 0.0481 0.0315 0.8463
HawkI-Real $(-20^{\circ},330^{\circ})$ Zero123++ 0.7132 25.5205 17.4653 0.0833 0.0977 0.7520
HawkI-Real $(-20^{\circ},330^{\circ})$ Stable Zero123 0.5730 23.9690 17.0343 0.0328 0.0726 0.7258
@@
Table 2: Quantitative Results. Evaluation on six metrics demonstrate the superior results of our method over prior work.
@@ 6 Conclusions, Limitations and Future Work
In this paper, we propose an approach that integrates the advantages of off-the-shelf 3D-based pretrained models within 3D-free paradigms for novel view synthesis, offering precise control over camera angle and elevation. Our method performs effectively on complex, in-the-wild images containing multiple objects and background information. We qualitatively and quantitatively demonstrate the benefits of our method over corresponding 3D and 3D-free baselines. One limitation of our method is its reliance on an inference-time optimization routine for each scene and viewpoint, which may hinder real-time performance. Achieving faster performance is a direction for future work. Additionally, extending our approach to NVS and 3D applications with real-world constraints (such as respecting contact points and relative sizes) for tasks like editing, object insertion, and composition presents promising directions for further research. Based on the current results, we also propose exploring the use of an image-conditioned model to achieve a higher level of view consistency as a future research direction.

### v4

@@ 5 Experiments and Results
Figure 7: Ablation Study on the use of regularization loss between angle embedding and optimized embedding In this experiment, we analyze the effect of adding a regularization term between the angle embedding ($e_{target}$) and the optimized embedding ($e_{view}$) on camera control results. The results show improvements in viewpoint consistency and style coherence when the regularization loss is applied.
@@
We utilize the HawkI-Syn Kothandaraman et al. (2023b) and HawkI-Real Kothandaraman et al. (2023b) datasets that feature complex scenes with multiple foreground objects and background. Both datasets provide images and text prompts to the model.
@@
We compare our method with state-of-the-art view synthesis methods: Zero123++ Shi et al. (2023a) and Stable Zero123 for 3D-based methods and HawkI Kothandaraman et al. (2023b) for 3D-free method.
@@ 5.1 Qualitative analysis
Our model is able to generate distinct viewpoints in the camera control to ensure consistency across generated views. Details are mentioned in the Zero123++ Shi et al. (2023a). We present qualitative representative results and comparisons with Zero123++ Shi et al. (2023a), HawkI Kothandaraman et al. (2023b), and StableZero123 at camera angles of $(30^{\circ},30^{\circ})$ and $(30^{\circ},270^{\circ})$ in Figures 5 and 6. Our method demonstrates superior scene reconstruction from all viewpoints compared to previous works. Specifically, results on HawkI-Syn in Figure 5 show that StableZero123 is largely ineffective. HawkI fails to capture the correct camera elevation in all cases except for the house image. While Zero123++ handles both elevation and azimuth, it struggles with background and detailed features. For instance, the pyramid in the first row lacks shadow information; the waterfall image in the second row appears unnatural; and the house in the last row blurs detailed features. Conversely, our model accurately reflects shadow characteristics in the pyramid, and reconstructs the details and background of the waterfall and house examples from various viewpoints.
@@ 5.2 Quantitative Evaluation
Following prior work Shi et al. (2023a); Kothandaraman et al. (2023b); Liu et al. (2023a), we evaluate our method using six metrics - (i) LPIPS Zhang et al. (2018): Quantifies the perceptual similarity between the generated and input images, with lower values indicating better performance. (ii) CLIP-Score Radford et al. (2021) and View-CLIP Score: Measure text-based alignment of the generated images. The CLIP score assesses alignment with both content and the $(\alpha_{elev},\alpha_{azi})$ viewpoint, while the View-CLIP Score focuses specifically on the viewpoint. Higher values are preferred. (iii) DINO Caron et al. (2021), SSCD Pizzi et al. (2022): DINO evaluates the semantic consistency of the generated images by comparing high-level feature embeddings extracted from a self-supervised vision transformer. DINO is trained not to ignore differences between subjects of the same class. Higher values indicate better preservation of semantic content across different views of the same scene. SSCD measures structural similarity between the generated images and their reference counterparts using learned feature representations. SSCD focuses on capturing fine-grained structural and contextual consistency. Higher values are preferred for better alignment with ground-truth structures. (iv) CLIP-I Ruiz et al. (2023): CLIP-I measures the cosine similarity between the embeddings of multi-view images and the input image within the CLIP space. (v) PSNR (Peak Signal-to-Noise Ratio) and SSIM Wang et al. (2004) (Structural Similarity Index): PSNR Quantifies the pixel-wise fidelity of the generated images relative to the reference images. PSNR is calculated as the logarithmic ratio of the maximum possible pixel value to the mean squared error between the two images. Higher values indicate better pixel-level accuracy. SSIM assesses perceptual similarity by comparing luminance, contrast, and structural information between the generated and reference images. SSIM is designed to measure structural consistency, with higher values reflecting closer perceptual alignment.
@@
Similar to the quantitative comparison performed by Zero123++, we use 10% of the overall data from the HawkI-Syn and HawkI-Real datasets as the validation set to compute the quantitative metrics. Table 1 and Table 5 shows that our model significantly outperforms the state-of-the-art across these evaluation metrics, reinforcing how our models stands out by incorporating the robust features of 3D-based NVS methods into a 3D-free optimization strategy, thereby capitalizing on the benefits of both approaches.
@@
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ DINO $\uparrow$ SSCD $\uparrow$ CLIP-I $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
HawkI-Syn $(30^{\circ},30^{\circ})$ Ours 0.5661 29.9563 0.4314 0.3638 0.8317 11.0664 0.3162
HawkI-Syn $(30^{\circ},30^{\circ})$ HawkI 0.5998 28.3786 0.3982 0.3519 0.8221 10.7092 0.2941
HawkI-Syn $(30^{\circ},30^{\circ})$ Zero123++ 0.5694 28.2555 0.4293 0.4605 0.8149 10.9923 0.3073
HawkI-Syn $(30^{\circ},30^{\circ})$ Stable Zero123 0.7178 21.3430 0.2108 0.2386 0.6467 9.2585 0.1954
HawkI-Syn $(30^{\circ},270^{\circ})$ Ours 0.5744 29.1800 0.4148 0.3684 0.8327 11.0661 0.3047
HawkI-Syn $(30^{\circ},270^{\circ})$ HawkI 0.5971 27.9540 0.3964 0.3473 0.8278 10.6303 0.2779
HawkI-Syn $(30^{\circ},270^{\circ})$ Zero123++ 0.6056 25.6665 0.2681 0.2195 0.7087 10.4395 0.2984
HawkI-Syn $(30^{\circ},270^{\circ})$ Stable Zero123 0.6785 23.1555 0.2119 0.2657 0.6456 9.4703 0.1673
HawkI-Real $(30^{\circ},30^{\circ})$ Ours 0.6201 29.8850 0.3346 0.2588 0.8152 9.4009 0.2184
HawkI-Real $(30^{\circ},30^{\circ})$ HawkI 0.6529 27.5847 0.2844 0.2269 0.7754 8.9257 0.2160
HawkI-Real $(30^{\circ},30^{\circ})$ Zero123++ 0.6253 27.9877 0.3315 0.3362 0.8023 9.2962 0.1990
HawkI-Real $(30^{\circ},30^{\circ})$ Stable Zero123 0.6614 23.0895 0.1781 0.1192 0.6569 7.7977 0.1684
HawkI-Real $(30^{\circ},270^{\circ})$ Ours 0.5868 30.5489 0.4126 0.3424 0.8708 10.6177 0.2687
HawkI-Real $(30^{\circ},270^{\circ})$ HawkI 0.6215 29.0488 0.3530 0.3363 0.8358 10.6472 0.2439
HawkI-Real $(30^{\circ},270^{\circ})$ Zero123++ 0.6302 27.5228 0.3145 0.2005 0.7529 9.8864 0.2484
HawkI-Real $(30^{\circ},270^{\circ})$ Stable Zero123 0.6268 21.1090 0.1750 0.0494 0.6500 8.3163 0.1637
@@
Table 1: Quantitative Results. Evaluation of seven metrics demonstrates the superior results of our method over prior work.
@@
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ DINO $\uparrow$ SSCD $\uparrow$ CLIP-I $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
HawkI-Syn $(30^{\circ},30^{\circ})$ w/ regularization 0.5661 29.9563 0.4314 0.3638 0.8317 11.0664 0.3162
HawkI-Syn $(30^{\circ},30^{\circ})$ w/o regularization 0.5867 28.5417 0.4122 0.3640 0.8243 10.8272 0.2954
HawkI-Real $(30^{\circ},30^{\circ})$ w/ regularization 0.6201 29.8850 0.3346 0.2588 0.8152 9.4009 0.2184
HawkI-Real $(30^{\circ},30^{\circ})$ w/o regularization 0.6257 29.0798 0.3357 0.2401 0.8231 9.1957 0.2014
HawkI-Syn $(30^{\circ},270^{\circ})$ w/ regularization 0.5744 29.1800 0.4148 0.3684 0.8327 11.0661 0.3047
HawkI-Syn $(30^{\circ},270^{\circ})$ w/o regularization 0.5952 28.9866 0.4098 0.3350 0.8248 10.8656 0.2850
HawkI-Real $(30^{\circ},270^{\circ})$ w/ regularization 0.5868 30.5489 0.4126 0.3424 0.8708 10.6177 0.2687
HawkI-Real $(30^{\circ},270^{\circ})$ w/o regularization 0.6114 29.9184 0.4003 0.3075 0.8541 10.2958 0.2615
HawkI-Syn $(-20^{\circ},210^{\circ})$ w/ regularization 0.5740 29.1144 0.4277 0.3529 0.8280 10.9697 0.2837
HawkI-Syn $(-20^{\circ},210^{\circ})$ w/o regularization 0.5860 28.6385 0.3969 0.3559 0.8171 10.8401 0.2792
HawkI-Real $(-20^{\circ},210^{\circ})$ w/ regularization 0.6185 30.6729 0.3610 0.2880 0.8448 10.3130 0.2223
HawkI-Real $(-20^{\circ},210^{\circ})$ w/o regularization 0.6338 29.1693 0.3817 0.2605 0.8263 10.0794 0.2117
HawkI-Syn $(-20^{\circ},330^{\circ})$ w/ regularization 0.5624 29.2144 0.4487 0.3892 0.8559 11.2175 0.3048
HawkI-Syn $(-20^{\circ},330^{\circ})$ w/o regularization 0.5714 28.4089 0.4476 0.3870 0.8492 11.0409 0.2947
HawkI-Real $(-20^{\circ},330^{\circ})$ w/ regularization 0.5925 29.5090 0.3899 0.3127 0.8689 10.6183 0.2971
HawkI-Real $(-20^{\circ},330^{\circ})$ w/o regularization 0.5894 28.8531 0.3704 0.2954 0.8506 10.5213 0.2828
@@
Table 2: Quantitative Results of Ablation Study. Evaluation of seven metrics demonstrates the superior results of the regularized method over the non-regularized one.
5.3 Ablation analysis: viewpoint regularization loss
To demonstrate the effectiveness of our approach in achieving camera control, we present results for scenes such as a Hawaiian beach and a waterfall. In both instances, the guidance images from Zero123++ fail to provide accurate direction to the model. For the Hawaiian beach scene, the output generated with the regularization term exhibits a more consistent style compared to the output produced without it. Despite the inaccuracies in the Zero123++ guidance images, the regularization term facilitates more reliable camera control than the results generated without it.
Similarly, in the waterfall scene, the regularization term enhances the consistency of the generated rock textures surrounding the waterfall. Without the regularization term, these textures are inconsistently represented; however, with it, the style is maintained more faithfully. Once again, Zero123++ does not provide accurate guidance in this case, underscoring the significant contribution of the regularization loss to improved control and visual coherence in the generated images. Detailed results from our ablation study are presented in Figure 7. Furthermore, the application of the regularization loss demonstrates performance improvements in quantitative evaluations, as shown in Table 2.
@@ 6 Conclusions, Limitations and Future Work
In this paper, we propose an approach that integrates the advantages of off-the-shelf 3D-based pretrained models within 3D-free paradigms for novel view synthesis, offering precise control over camera angle and elevation, without any additional 3D information. Our method performs effectively on complex, in-the-wild images containing multiple objects and background information. We qualitatively and quantitatively demonstrate the benefits of our method over corresponding 3D and 3D-free baselines. One limitation of our method is its reliance on an inference-time optimization routine for each scene and viewpoint, which may hinder real-time performance. Achieving faster performance is a direction for future work. Additionally, extending our approach to NVS and 3D applications with real-world constraints (such as respecting contact points and relative sizes) for tasks like editing, object insertion, and composition presents promising directions for further research. Based on the current results, we also propose exploring the use of an image-conditioned model to achieve a higher level of view consistency as a future research direction. As mentioned in the qualitative analysis, Due to its design, Zero123++ is limited to generating only distinct fixed views. While this approach improves consistency by leveraging Stable Diffusion’s priors, it restricts the model’s ability to generate views beyond these predefined angles, limiting flexibility in exploring arbitrary perspectives. Future work could explore enabling camera control from any angle, while addressing 3D prior model’s challenges like preserving source view attributes and mitigating issues from incorrect pose information to improve consistency and accuracy.
@@ A.1 Additional Results
A.2 Computation Time
Table 3 presents a comparison of memory consumption and computation time across state-of-the-art 3D-prior models, including Zero123++, Stable Zero123, and ZeroNVS. Among these models, Zero123++ demonstrates the shortest computation time, requiring only 20 seconds, while other methods are significantly slower.
Our approach utilizes Zero123++ for generating 3D prior information, ensuring that the computational cost remains minimal. Importantly, the generation of multi-view guidance images does not introduce any additional overhead, as this step is performed using the most computationally efficient model in this category. This demonstrates that our method is well-suited for scalable and real-time applications, maintaining efficiency while incorporating powerful 3D prior information.
Model Memory Consumption Computation Time
Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Stable Zero123 39.3 GB / 40.0 GB (37,479 MiB) 1,278 sec
ZeroNVS 33.48 GB / 40.0 GB (31,929 MiB) 7,500 sec
Table 3: Comparison of computation times for 3D-prior models. Among the prior works in NVS frequently mentioned, including Zero123++, Stable Zero123, and ZeroNVS, the Zero123++ model has the shortest computation time. Our research applies the Zero123++ model, which has the lowest computation time among 3D-prior models, to obtain 3D prior information without requiring any additional computation time.
Table 4 provides a detailed breakdown of memory usage and computation time for the optimization and inference steps in our method. The optimization process requires 387 seconds (Optimization 367 sec + Zero123++ 20 sec), and the inference step is highly efficient, taking only 6 seconds per image. Notably, the memory consumption remains consistent across optimization and inference, excluding the Zero123++ computation, comparable to other competitive methods.
This breakdown highlights that the inclusion of the Zero123++ step in our approach does not result in excessive computational time. Instead, our method achieves high-quality multi-view synthesis while maintaining practical memory and runtime efficiency. Furthermore, the results illustrate that our approach is capable of integrating multi-view guidance and reconstructing images with enhanced fidelity without compromising scalability or practicality.
This experiment was conducted using an A100 GPU (40.0 GB) for all models to ensure a fair comparison. The GPU memory consumption for each model is reported in the worst-case scenario, and the computation time is measured based on the time taken for the model to fully generate an image. The results in Table 3 and Table 4 emphasize that the proposed method effectively balances computational efficiency with enhanced performance. By leveraging Zero123++, the most efficient 3D-prior model, and incorporating lightweight optimization techniques, our approach ensures minimal computational costs while achieving significant improvements in output quality. These results validate the feasibility of the method for real-world applications, demonstrating both its scalability and practicality.
Model Step Memory Consumption Computation Time
HawkI Optimization 7.20 GB / 40.0 GB (6,875 MiB) 395 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 8.43 GB / 40.0 GB (8,045 MiB) 401 sec
Ours (w/o regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,879 MiB) 372 sec
Inference 8.43 GB / 40.0 GB (8,049 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 398 sec
Ours (w/ regloss) Zero123++ 10.18 GB / 40.0 GB (9,715 MiB) 20 sec
Optimization 7.21 GB / 40.0 GB (6,885 MiB) 367 sec
Inference 8.43 GB / 40.0 GB (8,045 MiB) 6 sec (each image)
Total 10.18 GB / 40.0 GB (9,715 MiB) 393 sec
Table 4: Detailed Step-wise Comparison. Even when applying Zero123++ to our methodology, the additional GPU memory consumption is relatively small, at 2.97GB (10.18GB - 7.21GB), and it takes only 20 seconds to generate the guidance image using Zero123++. From an overall perspective, HawkI takes 401 seconds to complete optimization and generate the first image through inference, while Ours (w/o regloss) takes 398 seconds, and Ours (w/ regloss) takes 393 seconds. This demonstrates that our methodology does not result in significant differences in computation time or memory consumption while achieving better performance compared to existing methods. Total memory consumption refers to the worst case, the computation time indicates the total execution time. i.e., the time taken for the model to run and output the first image.
A.3 Quantitative Evaluation Results
Dataset, Angle, Method LPIPS $\downarrow$ CLIP $\uparrow$ DINO $\uparrow$ SSCD $\uparrow$ CLIP-I $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$
HawkI-Syn $(-20^{\circ},210^{\circ})$ Ours 0.5740 29.1144 0.4277 0.3529 0.8280 10.9697 0.2837
HawkI-Syn $(-20^{\circ},210^{\circ})$ HawkI 0.6024 27.7407 0.3831 0.3494 0.8226 10.5667 0.2744
HawkI-Syn $(-20^{\circ},210^{\circ})$ Zero123++ 0.6037 24.4148 0.2936 0.3021 0.7309 10.7458 0.2803
HawkI-Syn $(-20^{\circ},210^{\circ})$ Stable Zero123 0.7452 20.7860 0.0852 0.0996 0.5634 6.3887 0.0971
HawkI-Syn $(20^{\circ},330^{\circ})$ Ours 0.5624 29.2144 0.4487 0.3892 0.8559 11.2175 0.3048
HawkI-Syn $(20^{\circ},330^{\circ})$ HawkI 0.5943 27.5738 0.4080 0.3532 0.8152 10.8882 0.2759
HawkI-Syn $(20^{\circ},330^{\circ})$ Zero123++ 0.5652 25.8831 0.4305 0.4431 0.7932 11.1130 0.2936
HawkI-Syn $(20^{\circ},330^{\circ})$ Stable Zero123 0.6332 23.2087 0.3366 0.3393 0.6890 9.1852 0.1943
HawkI-Real $(-20^{\circ},210^{\circ})$ Ours 0.6185 30.6729 0.3610 0.2880 0.8448 10.3130 0.2223
HawkI-Real $(-20^{\circ},210^{\circ})$ HawkI 0.6464 28.7500 0.3567 0.2697 0.8001 9.6859 0.2145
HawkI-Real $(-20^{\circ},210^{\circ})$ Zero123++ 0.6816 24.7083 0.2101 0.1706 0.6434 8.6865 0.2194
HawkI-Real $(-20^{\circ},210^{\circ})$ Stable Zero123 0.6650 21.5791 0.1564 0.0225 0.5850 7.4097 0.1681
HawkI-Real $(-20^{\circ},330^{\circ})$ Ours 0.5925 29.5090 0.3899 0.3127 0.8689 10.6183 0.2971
HawkI-Real $(-20^{\circ},330^{\circ})$ HawkI 0.6283 27.5200 0.3228 0.2406 0.8383 10.4706 0.2787
HawkI-Real $(-20^{\circ},330^{\circ})$ Zero123++ 0.5978 26.1550 0.3735 0.3080 0.8043 10.5917 0.2953
HawkI-Real $(-20^{\circ},330^{\circ})$ Stable Zero123 0.6673 25.6611 0.2667 0.1998 0.7249 9.0786 0.1653
Table 5: Quantitative Results. Evaluation of seven metrics demonstrates the superior results of our method over prior work.